OpenStudy (anonymous):

Solve equation with parameter k | x + 5 - k | = | x - 2 |

OpenStudy (anonymous):

solve \[x+5-k=x-2\] and also \[x+5-k=2-x\]

OpenStudy (anonymous):

OK. thanks